OpenStudy (jollypikachu300):

WILL FAN AND MEDAL NEED HELP Solve for n. A.3/4 B.4/3 C.1 3/4 D.1 4/3

HanAkoSolo (jamierox4ev3r):

Hi, what is the equation?

OpenStudy (jollypikachu300):

its fraction

OpenStudy (jollypikachu300):

thats the question

OpenStudy (jollypikachu300):

and the anwser

OpenStudy (anonymous):

obviously cats lifting weights uhm

HanAkoSolo (jamierox4ev3r):

no i mean...you said to solve for n. That means that there must be an equation of sorts. I have no clue what I'm solving for

OpenStudy (jollypikachu300):

ok sorry here is the question

OpenStudy (jollypikachu300):

solve for n 3/4 times n equals 1

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }n = 1\] ?

OpenStudy (anonymous):

\[\frac{ 3 n}{ 4 } = 1\]

OpenStudy (jollypikachu300):

yes

OpenStudy (anonymous):

solve for n.

OpenStudy (jollypikachu300):

i suck at fraction

OpenStudy (anonymous):

multiply 4 to the other side.

HanAkoSolo (jamierox4ev3r):

a simple way to do this is to multiply 4 to both sides

OpenStudy (anonymous):

\[3n = 1*4\] now there is no fraction. Solve for n

HanAkoSolo (jamierox4ev3r):

oops. lol. you're fast

OpenStudy (anonymous):

cat instincts. . .

OpenStudy (jollypikachu300):

lol